### v8

# Quantum chaos and the arrow of time

Abstract
Classical physics provides an arrow of time in the form of the second law of thermodynamics. But a clear picture of the quantum origin of the arrow of time has been lacking so far. In this letter, we show that an arrow of time arises in quantum chaotic systems. We show that, for an isolated quantum system which is also chaotic, the change in entropy is non-negative when the system is perturbed. We show that this gives rise to a new profound constraint on the off-diagonal terms in eigenstate thermalization hypothesis (ETH) statement. In case of an integrable system, the second law does not hold true because the system does not thermalize to a generalized Gibbs ensemble after a finite perturbation.
All physical systems around us become more and more disordered as time progresses unless an outside agent spends energy and record information for the upkeep of the system. This phenomenon has a beautiful statistical reasoning. There simply are too many possible disordered states compare to the number of possible ordered states. So, the system ends up most likely in a disordered state [1]. Entropy is the quantity which measures the extent of disorder. In terms of entropy, this general observation gives rise to the second law of thermodynamics. It states that the entropy of an isolated system does not decrease. This is Planck’s statement of the second law of thermodynamics [2]. It gives rise to an arrow of time (called Time’s arrow in [3]). These results are familiar in the realm of classical physics. But our physical world is governed by quantum mechanics at microscopic level and so far a complete understanding of the quantum version of the second law is lacking.
In this letter, we show that an arrow of time arises naturally in an isolated quantum system which is chaotic. Quantum chaos in many-body systems and the related question of thermalization of these systems have been areas of intense research in the last few decades [4]. Traditionally, quantum chaos is identified from the study of energy level statistics [5], while thermalization in a quantum system follows from eigenstate thermalization hypothesis (ETH) [6, 7]. The main idea of ETH comes from Berry’s conjecture [8] which states that eigenstates of quantum chaotic systems are thermal in an observational sense. ETH explains how observables thermalize in a quantum system. So, it is the route to statistical mechanics for quantum systems. ETH has been shown to hold true for physically sensible observables of quantum chaotic systems while it does not hold true for quantum integrable systems. So for the rest of the letter, we will assume that quantum chaos and ETH imply each other. In technical aspect, we found a new profound constraint on the ETH statement.
It has previously been shown that the change in the diagonal entropy is non-negative under doubly stochastic evolution starting from a passive density matrix [9]. This follows from the convex nature of the function $x\log x$. This result is true even for integrable systems. But integrable systems do not thermalize after a finite perturbation. In the present work, we will be working with pure states which is the most appropriate approach for an isolated system. In the spirit of Berry’s conjecture and ETH, we will be working with individual energy eigenstates.
Just as we cannot keep track of the positions and momenta of all gas molecules in the classical theory of gas, we cannot keep track of the complex coefficients of the exponentially large number of energy eigenstates of a quantum system. So in quantum mechanics, we keep track of only certain observables which can be easily measured or calculated. These can be local operators or global operators but they are few-body operators. If we observe that these observables thermalize then we say that the system thermalizes. ETH is a criteria for thermalization of such observables so that their long time expectation values can be described by quantum statistical mechanics. Inevitably, the quantum operators corresponding to these observables are hermitian and non-fermionic in nature.
Consider an operator $\mathcal{O}$ which corresponds to one such observable. ETH states that the matrix elements $O_{mn}$ in energy eigenbasis are of the form
$\langle m|\mathcal{O}|n\rangle\equiv O_{mn}=O(\bar{E})\delta_{mn}+e^{-S(\bar{E})/2}f(\bar{E},\omega)R_{mn}$ (1)
where $|m\rangle$ and $|n\rangle$ are energy eigenstates with energies $E_{m}$ and $E_{n}$ respectively and $\bar{E}=(E_{m}+E_{n})/2$, $\omega=E_{m}-E_{n}$. $S(\bar{E})$ is the entropy at energy $\bar{E}$. $\mathcal{O}(\bar{E})$ is equal to the expectation value of $\mathcal{O}$ in the microcanonical ensemble at energy $\bar{E}$ or other ensembles by the equivalence of ensembles. $\mathcal{O}(\bar{E})$ and $f(\bar{E},\omega)$ are smooth functions of their arguments. $f(\bar{E},\omega)$ also falls exponentially for large $|\omega|$ [10]. $R_{mn}$ are pseudo-random variables with zero mean and unit variance. $\mathcal{O}$ is hermitian and hence
$O_{nm}=O^{*}_{mn},\;R_{nm}=R^{*}_{mn},\;f(\bar{E},-\omega)=f^{*}(\bar{E},\omega)$ (2)
In this work, we found that the second law of thermodynamics enforces that
$|f(\bar{E},\omega)|$ is a monotonically increasing function of $S(\bar{E})$.
The $\bar{E}$ dependence of $f(\bar{E},\omega)$ has never been studied in detail in previous studies of ETH and quantum chaos. $\bar{E}$ is usually used only to fix $S(\bar{E})$ and the effective temperature.
The arrow of time has two different manifestations. The first manifestation is thermalization of non-equilibrium excited states which are typical [11]. If the statement of ETH is satisfied then the system thermalizes to the thermal state specified by the fixed conserved charges in the long time limit. This scenario is most popularly studied as quantum quenches where the non-equilibrium state is prepared by changing the Hamiltonian of the system [12, 13]. The initial Hamiltonian and the final Hamiltonian are different.
One subtlety when preparing states using quantum quenches is that even integrable systems thermalize to generalized Gibbs ensembles (GGEs) in the large system size limit, if the initial state has no long range correlation. Thermalization of quantum integrable systems after quantum quenches has been a topic of great interest in the last two decades [14, 15, 16, 12, 13]. The non-equilibrium states are generalized typical states taking into account all the other conserved charges of the integrable system. This result simply follows from central limit theorem and the local nature of the Hamiltonians [17]. The difference between chaotic systems and integrable systems arises when we consider finite perturbations for a finite duration of time (also called bump quenches in [18] and critical to critical quench in [16]). The initial Hamiltonian and the final Hamiltonian are same. A chaotic system still thermalizes after such finite perturbations [18] while integrable systems do not thermalize [16]. The effective temperature and other chemical potentials become imaginary. So in this manner, finite perturbations or bump quenches bring out the stark contrast between non-equilibrium dynamics of quantum chaotic systems and quantum integrable systems.
The second manifestation of the arrow of time is the quantum version of the second law of thermodynamics. Consider a quantum chaotic system which is initially in a thermal state (a pure state but thermalized). If we perturb the system, the system would be taken out of equilibrium but eventually the system would return back to thermal equilibrium. The second law would imply that the change in entropy ($\Delta S$) is non-negative. We will show that this is indeed the case for a quantum chaotic system. We show this by keeping track of the change in total energy of the system ($\Delta E$). The change in energy is positive if the initial temperature of the system is positive. If the system was initially at infinite temperature then the change in the energy is zero.
$\Delta E=T\Delta S$ (3)
where $\Delta E$ and $\Delta S$ are changes in energy and entropy. We have set the Planck constant $\hbar=1$ and the Boltzmann constant $k_{B}=1$. For systems with finite dimensional Fock space, if the initial temperature is negative, the change in energy is negative and the change in entropy is positive again. Figure 1 is the schematic diagram of this result. One can also consider changes in other conserved charges by considering the change in the expectation value of an effective Hamiltonian which include the conserved charges with appropriate values of the chemical potentials. Our result in terms of the change in energy is the Kelvin’s form of the second law of thermodynamics [19].
Figure 1: The density of states ($e^{S}$) versus energy ($E$) plot for the chaotic XXZ spin chain. $T_{eff}$ is effective temperature if we are working with energy eigenstates. Perturbation of the system from a thermalized state always leads to a non-negative change in the entropy $\Delta S\geq 0$.
Consider a system with the time-independent Hamiltonian $H_{0}$ and initially in a state $|\psi\rangle$. The system is perturbed by a time-dependent term $\lambda(t)\mathcal{O}$ where $\mathcal{O}$ is a non-fermionic, hermitian operator. The total Hamiltonian is
$H(t)=H_{0}+\lambda(t)\mathcal{O}$ (4)
The perturbation is for a finite duration of time, say, from time $-d$ to $d$. $\lambda(t)$ is real and can be non-zero only for $-d<t<d$. This perturbing term is not necessarily small but it also cannot be arbitrarily big. The change in energy of the system is
$\Delta E=\langle\psi(t_{f})|H_{0}|\psi(t_{f})\rangle-\langle\psi(t_{i})|H_{0}|\psi(t_{i})\rangle$ (5)
where $t_{i}<-d,d<t_{f}$. Our main goal is to show that $\Delta E$ is positive (negative, zero) if the initial temperature is positive (negative, infinite) when $H_{0}$ is chaotic and $\mathcal{O}$ satisfies ETH. This would imply that $\Delta S\geq 0$. The initial state is taken to be an energy eigenstate $|n\rangle$. The entropy is most easily calculated from the effective temperature which in turn is calculated using the equivalence of ensembles.
Using the Dyson series expansion of the time evolution operators, we can perform an expansion of $\Delta E$ in powers of $\lambda(t)$. We will concentrate on the leading term which is given by
$\Delta E=\sum_{m}\Delta E_{m}=\sum_{m}(E_{m}-E_{n})|\tilde{\lambda}(E_{m}-E_{n})|^{2}|O_{mn}|^{2}$ (6)
[…]
where $A(\omega)=-2\,\text{Im}\;G_{R}(\omega)$ is the spectral function. $G_{R}(t,t^{\prime})=-i\theta(t-t^{\prime})\langle[\mathcal{O}_{I}(t),\mathcal{O}_{I}(t^{\prime})]\rangle$ is the retarded Green’s function where $\mathcal{O}_{I}(t)$ is the perturbing operator in the interaction picture. For a non-fermionic operator, the above expression is positive (negative or zero) in a thermal state of positive (negative or infinite) temperature. Invoking Berry’s conjecture and ETH, we propose that the leading term in (6) and the thermal value should match except for exponentially suppressed corrections. This would not hold true for integrable systems.
We consider the XXZ spin chains with open boundary condition. With a large next-to-nearest neighbour interaction, the system is chaotic [20]. We consider system of size $L$ with the number of up-spins $N$. The system is perturbed using the global spin-current operator which preserves the total magnetization.
$H_{0}=\sum_{i=1}^{L-1}\left[J_{xy}\left(S_{i}^{x}S_{i+1}^{x}+S_{i}^{y}S_{i+1}^{y}\right)+J_{z}S_{i}^{z}S_{i+1}^{z}\right]$
$+\sum_{i=1}^{L-2}J^{\prime}_{z}S_{i}^{z}S_{i+2}^{z}$ (8)
$\mathcal{O}=\sum_{i=1}^{L-1}J_{xy}\left(S_{i}^{x}S_{i+1}^{x}-S_{i}^{y}S_{i+1}^{y}\right)$ (9)
We take $J_{xy}=1$ and $J_{z}=J^{\prime}_{z}=0.5$  [^1] The Mathematica notebook including the codes and the results is available as an ancillary file at https://arxiv.org/abs/2212.03914. Figure 2 are the plots of the leading term in $\Delta E$ after a unit delta function perturbation $\lambda(t)=\delta(t)$ in each of the energy eigenstates. In the chaotic system, the leading term starting from different eigenstates match the thermal values considering both canonical and microcanonical ensembles. As one expects, the matching is better for larger systems. It is remarkable that the eigenstates of the chaotic $L=12$ system still shows thermal behaviour for the leading term of $\Delta E$ even though the dimension of the Fock space is only 792. For thermalization studies of chaotic systems, one usually need Fock space with dimensions of the order of $30000$ [22, 23]. In case of the integrable system, the leading term starting from different eigenstates wildly vary compare to the thermal values.
Figure 2: The leading term of $\Delta E$ for chaotic and non-chaotic XXZ spin chains after a unit delta function perturbation $\lambda(t)=\delta(t)$. The blue dots are for the different eigenstates of energy $E_{n}$ as the initial state, the yellow plots are for thermal states with energy $E_{n}$ and the green dots are for microcanical ensembles.
The left panel in Figure 3 is the plot of $\Delta E_{m}$ as a function of $E_{m}$ for a fixed $E_{n}$ with effective temperature $T_{eff}=2$. We can see the exponential fall-off of $f(\bar{E},\omega)$ for large $|\omega|$. We can also see the role of the psuedo-random elements $R_{mn}$, albeit without the signs. The right panel in Figure 3 is the plot of the sum of $\Delta E_{m}$ from $-\omega$ to $\omega$ as a function of $\omega$. It shows that the bulk of $\Delta E$ value comes from the small $\omega$ region.
Figure 3: Left: $\Delta E_{m}$ as a function of $E_{m}$. Right: Sum of $\Delta E_{m}$ from $-\omega$ to $\omega$ as a function of $\omega$ for different eigenstates $|n\rangle$ with effective temperatures $T_{eff}=1,2,3$.
The naive substitution of ETH in (6) goes wrong in the details involving the integration contour in the $\bar{E}-\omega$ plane. One has to be careful about the $\bar{E}$ variation. The top-left panel in Figure 4 is an illustration of the $\bar{E}-\omega$ plane. The direction of $\omega$ and $\bar{E}$ axes are along the diagonals. The red line is the integration contour where we have consider only the values close to the $\bar{E}$ axis since $f(\bar{E},\omega)$ is exponentially suppressed for large $|\omega|$. Using the hermitian nature of $f(\bar{E},\omega)$, the right half of the contour can be rotated by $90^{\circ}$ clockwise. The integration along the two segments will cancel if not for the $\bar{E}$ variation which is represented by the blue contour. So, the $\bar{E}$ dependence of the integrand makes the total integral match with the thermal expectation value. This suggests that $|f(\bar{E},\omega)|$ is a monotonically increasing function of $S(\bar{E})$.
We calculate $|f(\bar{E},\omega)|$ as a function of $\bar{E}$ for the global spin-current operator in the chaotic XXZ spin chain. The three plots of Figure 4 are plots of $|f(\bar{E},\omega)|$ as a function of the effective inverse temperature of $\bar{E}$. These plots show that $|f(\bar{E},\omega)|$ is indeed a monotonically increasing function of $S(\bar{E})$ or the magnitude of the effective temperature $|T_{eff}|$. In all reasonable physical systems, the entropy is a monotonically increasing function of $|T_{eff}|$.
The edges in the energy spectrum of the XXZ spin chain are less chaotic [20]. This results in the unexpected behaviour at low temperature in the plots of $|f(\bar{E},\omega)|$. To verify this, we also calculated $|f(\bar{E},\omega)|$ in the highly chaotic $(q=4)$ SYK model [24, 25]. It is well known that even the edges of the spectrum are chaotic in SYK model [26]. We consider the occupation number at a single site as the observable so it corresponds to a local operator. Figure 5 are the plots of $|f(\bar{E},\omega)|$ as a function of the effective inverse temperature of $\bar{E}$ [21]. In this case, we find the expected behaviour even at very low temperature.
We also calculated the higher order terms of $\lambda(t)$ in $\Delta E$. For reasonable perturbation strength, the higher order terms are small compare to the leading term. $R_{mn}$ are not completely random [27]. If they were completely random, a full analytic proof of non-negative $\Delta S$ is possible. The odd power terms of $\lambda(t)$ are vanishing unless $\lambda(t)$ has a large odd function component or $\mathcal{O}$ is not time reversal invariant. The higher order terms can be used to constrain the higher moments of $R_{mn}$.
We also studied perturbations with large values of $\lambda(t)$ numerically. $\Delta E$ remains always positive (negative) for positive (negative) initial temperature. These large perturbations are not relevant for our discussion of the arrow of time.
Figure 4: Top-left: $\bar{E}-\omega$ plane and the contour of the integration. Other panels: $|f(\bar{E},\omega)|$ as a function of the effective inverse temperature $1/T_{eff}(\bar{E})$ for different $|\omega|$. The observable is the total spin current in the chaotic XXZ spin chain.
Figure 5: $|f(\bar{E},\omega)|$ as a function of the effective inverse temperature $1/T_{eff}(\bar{E})$ for different $|\omega|$. The observable is the occupation number at a single site in the chaotic $(q=4)$ SYK model.
We can still reduced the entropy by using a fine-tuned perturbation. This can be accomplished by predominantly turning on modes of $\tilde{\lambda}(\omega)$ which allows transition to only lower energy levels. But this fine-tuning requires the knowledge of a large part of the spectrum and precise knowledge of the initial state. This is reminiscent of the Maxwell’s demon problem in classical thermodynamics. In the present case, finding the energy levels at the required precision involves performing highly precise measurements which increase the total entropy of the system and the outside agent.
Summary. We show that quantum chaotic systems inherently possess an arrow of time. Working at leading order in perturbation theory, we show that the second law of thermodynamics constrains the off-diagonal terms in ETH statement. The new constraint has to be taken into account when one is working beyond probe limit. It would be worthwhile to study the higher order terms in perturbation theory and see how they constrain higher point functions.
Acknowledgements.
The author would like to thank Gautam Mandal for helpful discussions during one of which the main idea for this paper arose. The author is fully supported by the Department of Science and Technology (Government of India) under the INSPIRE Faculty fellowship scheme (Ref. No. DST/INSPIRE/04/2020/002105).